OpenStudy (anonymous):

Which does *not* show a direct variation between x and y? A. y=4/x B.y= 0.5x C. y=7x D.y= x/9 Im thinking it should be "B" am I right?

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

hmmm

OpenStudy (alexandervonhumboldt2):

oh wait.0 B is not correct for sure.

OpenStudy (alexandervonhumboldt2):

y=0.5x k=0.5 satisfies y=kx form. so it is direct

OpenStudy (anonymous):

okay hmm

OpenStudy (anonymous):

What about 7x its a whole number?

OpenStudy (alexandervonhumboldt2):

answer is not corerect, but 7 is a whole numbner

OpenStudy (anonymous):

What about D?

OpenStudy (alexandervonhumboldt2):

no. k=1/9 in this case.

OpenStudy (anonymous):

Wait so how would it be A then

OpenStudy (alexandervonhumboldt2):

yes it is A. it has form y=k/x not y=kx

OpenStudy (anonymous):

Because it fliped ahhh

OpenStudy (anonymous):

^*it is

OpenStudy (anonymous):

Thanks @AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

basically, a direct variation passes though 0. y=4/x is a hyperbola, which cannot be in 0.

OpenStudy (nincompoop):

definition definition and definition

OpenStudy (alexandervonhumboldt2):

i mean orgin

OpenStudy (nincompoop):

what does it mean by direct variation between one aspect and another aspect